OpenStudy (anonymous):

can you check this for me?!?! Leander rolls a standard die twice. Which of the following compound events has the greatest chance of occurring? the probability of 2 and 2 the probability of 2 and 5 the probability of 5 and even the probability of even and even

OpenStudy (anonymous):

it would be (even and even) correct?

OpenStudy (anonymous):

the probability of 2 and 2 is 1/6*1/6=1/36 this is the same for 2 and 5 as the probability of getting a 2 is equal as the probability of getting a 5 the probability of 5 and even is 1/6*1/2=1/12 since there is one 5 and there are 3 evens in a sample space of 6 the probability of even and even is therefore 1/2*1/2=1/4 which has the greatest probability of occurring

OpenStudy (anonymous):

YAY! i was right!